Calculate work done by a man of mass 60kg who climbs up a stair consisting of 15 step.Given of each steps is 30cm.
1
Expert's answer
2020-11-03T10:32:07-0500

By the definition of the work we have:


"W=Fh=mgh=60\ kg\cdot 9.8\ \dfrac{m}{s^2}\cdot 0.3\ m\cdot 15=2646\ J."

Answer:

"W=2646\ J."